OpenStudy (fanduekisses):

CALCULUS DERIVATIVES: Air is being released from a spherical balloon at the rate of 3 cu. in./min. What is the rate of change of the radius of the balloon when the radius of the balloon is 2 in.?

OpenStudy (fanduekisses):

So, I need to find dr/dt right?

OpenStudy (perl):

V = 4pi/3*r^3 dV/dt = dV/dr * dr/dt dV/dt = 4pi/3 * 3 * r^2 * dr/dt

OpenStudy (fanduekisses):

why v=4pi/3&r^3?

OpenStudy (fanduekisses):

Doesn't v= 4/3 *pi *r^3 ?

OpenStudy (perl):

thats the same thing

OpenStudy (perl):

4pi/3 = 4/3 * pi

OpenStudy (fanduekisses):

ohh yeah ok

OpenStudy (perl):



OpenStudy (fanduekisses):

so dr/dt= 3/16pi ?

OpenStudy (perl):

we have to be careful

OpenStudy (fanduekisses):

We have to rationalize?

OpenStudy (perl):

my draw button is frozen :/

OpenStudy (perl):

ok did you get dV/dt = 4pi* r^2 * dr/dt

OpenStudy (perl):

Air is being released from a spherical balloon at the rate of 3 cu. in./min. What is the rate of change of the radius of the balloon when the radius of the balloon is 2 in.? dV/dt = -3 r= 2 -3 = 4pi* (2)^2 * dr/dt

OpenStudy (fanduekisses):

so the 3's cancel out

OpenStudy (perl):

right

OpenStudy (fanduekisses):

ok